What is the radius of x^2 -y^2=4
Found 3 solutions by addingup, ikleyn, MathTherapy:
Answer by addingup(3677)

x^2 - y^2 = 4
The equation is;
x2 + y2 = r2
take the square root on all sides:
x - y = 2 this is your radius

Answer by ikleyn(52781)

.
The equation is NOT an equation of a circle.

It is an equation of a hyperbola.

Therefore, in your post the question is not relevant to the given part,
and, in opposite, the given part is not relevant to the question.


The post is defective.



Answer by MathTherapy(10552)

What is the radius of x^2 -y^2=4
HYPERBOLA.